[Q1] Suppose that PAlB) 04 and P(B0.5. Determine the following (a) P(AnB) (b) PAnB)
Answer #1

Solution:

Given that,

P ( A / B ) = P (A )

P ( A ) = 0.4

P ( B ) = 0.5

a) P ( A   \cap  B ) = P ( A ) * P ( B )

= 0.4 * 0.5

= 0.20

P ( A   \cap  B ) = 0.20

b )   P ( A "  \cap  B ) = P ( B ) - P (A )

= 0.5 - 0.4

= 0.1

P ( A "  \cap  B ) = 0.1